# New Bounds on Quotient Polynomials with Applications to Exact Divisibility and Divisibility Testing of Sparse Polynomials

### v3

@@ Abstract
We prove that for polynomials $f,g\in\mathbb{C}[x]$ such that $g$ divides $f$, the $\ell_{2}$-norm of the quotient polynomial $f/g$ is bounded by $\left\|{f/g}\right\|_{2}\leq\left\|{f}\right\|_{1}\cdot\tilde{O}(\left\|{g}\right\|_{0}^{3}\deg^{2}{f})^{\left\|{g}\right\|_{0}-1}$, where $\|g\|_{0}$ is the sparsity of $g$ (the number of its non-zero coefficients). This improves upon the previously known exponential (in $\deg(f)$) bounds for general polynomials. Our result implies that the trivial long division algorithm runs in quasi-linear time relative to the input size and number of terms of the quotient polynomial $f/g$, thus solving a long-standing problem on exact divisibility of sparse polynomials. Prior to our work it was not even known that the representation size of the quotient polynomial is sub-quadratic in its number of terms, i.e., that it is sub-quadratic in $\deg{f}$.
We also study the problem of bounding the number of terms of $f/g$ in some special cases. When $f,g\in\mathbb{Z}[x]$ and $g$ is a cyclotomic-free (i.e., it has no cyclotomic factors) trinomial, we prove that $\left\|{f/g}\right\|_{0}\leq O(\left\|{f}\right\|_{0}\text{size}\left(f\right)^{2}\cdot\log^{6}{\deg{g}})$. When $g$ is a cyclotomic-free binomial, we prove that the sparsity is at most $O(\left\|{f}\right\|_{0}(\log{\left\|{f}\right\|_{0}}+\log{\left\|{f}\right\|_{\infty}}))$. Both upper bounds are polynomial in the input-size. Without assuming cyclotomic-freeness, there are simple examples that show that both sparsities can be exponentially large. Leveraging these results, we provide a polynomial-time algorithm for deciding whether a cyclotomic-free trinomial divides a sparse polynomial over the integers. In addition, we show that testing divisibility by a binomial can be done in $\tilde{O}(n^{2})$ time over $\mathbb{Z}$ and over finite fields.
As our last result, we present a polynomial time algorithm for testing divisibility by pentanomials over small finite fields when $\deg{f}=\tilde{O}(\deg{g})$.
All our algorithms are deterministic.
@@ 1 Introduction
where $\left\|{f}\right\|_{0}$ and $\left\|{f}\right\|_{\infty}$ are the number of non-zero coefficients in $f$, and the maximal coefficient of $f$ (in absolute value), respectively.111 As we think of $\left\|{f}\right\|_{0}$ as being considerably smaller than $\log_{2}{\left\|{f}\right\|_{\infty}}$ and $\log_{2}{\deg{f}}$, we do not pay attention to the case where bit lengths of coefficients vary. Similarly, over a finite field $\mathbb{F}$ the representation size is
$\text{size}\left(f\right)=\left\|{f}\right\|_{0}(\log_{2}{\left|\mathbb{F}\right|}+\log_{2}{\deg{f}}).$
This makes the sparse representation particularly relevant to applications in which the polynomial has a large degree and most of its coefficients are zero. In particular, it is used by computer algebra systems and libraries such as Maple [MP13], Mathematica [Wol99], Sage [S+08], and Singular [Sch03].
@@
One of the long-standing open problems concerns exact divisibility of sparse polynomials. There are two versions to this problem. The decision problem asks, given two sparse polynomials $f$ and $g$, to decide whether $g$ divides $f$. The search problem, or the exact division problem, asks to compute the quotient (also called co-factor) polynomial $f/g$ when it is known that $g$ divides $f$. The surveys of Davenport and Carette [DC09] and of Roche [Roc18] thoroughly discuss these problems.
The two problems are closely related but also differ in what we can hope to achieve. In the decision problem, the output is a single bit, allowing us to analyze the complexity in terms of the input size. Conversely, in the search problem, the output size depends on the number of monomials in $f/g$, which may require an exponential number of bits to represent. Next, we will discuss each of these problems separately.
1.1 Exact Division
There are two difficulties in the exact division problem. The first stems from the fact that $f/g$ can have exponentially (in the input-size) many monomials. The second concerns the height (i.e. the maximum among the absolute values of the coefficients) of the quotient polynomial, when working over the integers. Prior to this work the only bounds on the size of its coefficients were exponential in $\deg{f}$ (see [Mig12]). [^2] These results are tight when $g$ is not a sparse polynomial. In particular, Davenport and Carette note that even in the case where $g$ exactly divides $f$, we need to perform $\Omega(\|f/g\|_{0}\cdot\|g\|_{0})$ operations on coefficients, and the number of bits required to represent a coefficient may be as large as $\deg f$ [DC09, Section III.B]. Therefore, in terms of the input size and $\|f/g\|_{0}$ the only provable bounds on the running time of the exact division algorithm were of the form
@@ 1.1 Exact Division
1.2 Divisibility Testing
Recall that in the divisibility testing problem the goal is to determine whether $g$ divides $f$ without having to compute the quotient and remainder. Thus, we seek a polynomial-time algorithm in the size of the input. It is easy to see that when $\deg{g}$ is small (or, similarly, $\deg{f}-\deg{g}$ is small), or when the quotient polynomial is sparse, the problem becomes easy.
However, in the general case, it is only known that, assuming the Extended Riemann Hypothesis (ERH), the divisibility problem is in coNP [GKO92]. This led Davenport and Carette to pose the following challenge:
Open Problem 1 (Challenge 3 of [DC09]).
Find a class of instances for which the question "does g divide f?" is NP-complete, or find an algorithm for the divisibility of polynomials which is polynomial-time. Failing this - find an algorithm for the divisibility of cyclotomic-free polynomials which is polynomial-time.
@@ 1.2 Divisibility Testing
In fact, even the next problem, that was stated in [Roc18] following [DC09], is still open. [^3] As usual, $\tilde{O}(t)=O\left(t\cdot\log^{O(1)}t\right)$ for a running time function $t$, and when no basis is mentioned then the log is taken in the natural base $e$
Open Problem 2 ([Roc18, Open Problem 3]).
Given two sparse polynomials $f,g\in R[x]$, develop an algorithm which determines whether $g$ divides $f$ exactly, using $\tilde{O}(T\log{\deg{f}})$ ring and bit operations, where $T$ is the total number of terms of $f$, $g$, $(f\text{ quo }g)$ and $(f\text{ rem }g)$ (the quotient and remainder).
@@
Note that 2 asks for an algorithm whose running time may be exponential in the input size.
There are several reasons to believe that divisibility testing is a hard problem. In [Pla84], Plaisted studied the problem of divisibility testing of sparse polynomials and proved that determining whether the quotient polynomial has a non-zero constant term is NP-hard and that determining the degree of the remainder polynomial is also NP-hard. Divisibility testing of sparse polynomials is a special case of a more general problem of divisibility testing of arithmetic circuits in one variable (of arbitrary degree). This problem too was shown to be NP-hard - as mentioend, testing whether $x^{n}-1$ divides a product of sparse polynomials [Pla77] is NP-Hard. While these results do not prove that the divisibility problem is NP-hard, they show that closely related problems are difficult to solve.
Davenport and Carette highlighted polynomials with cyclotomic factors in 1 because, since Plaisted’s work [Pla77], all NP-hard problems related to sparse polynomials involve cyclotomic polynomials. For example, many problems are known to be NP-hard when the polynomials involved can have cyclotomic factors, such as deciding whether there exists a non-trivial common divisor of two polynomials, testing whether a polynomial has a cyclotomic factor [Pla84], and determining if a polynomial is square-free [KS99]. Notably, it remains an open question whether these problems, and the NP-hard problem mentioned above, remain NP-hard when considering cyclotomic-free polynomials. For further discussion on the role of cyclotomic polynomials and the difficulties they pose, see the surveys [DC09, Roc18].
There are multiple obstacles hindering natural approaches, including long-division, computing the remainder, or testing evaluations, from efficiently solving the divisibility testing problem. We first note that the quotient of two sparse polynomials may not be sparse. For example, consider the equation
$\frac{x^{n}-1}{x-1}=\sum_{0\leq i<n}{x^{i}}.$
Here the size of the input is $O(\log n)$ whereas the size of the output is $\Omega(n)$. This rules out the simple long division algorithm and requires any solution to decide whether such a quotient exists without computing it explicitly. Similarly, we note that the remainder of the division of two sparse polynomials may not be sparse as well, and, moreover, may contain doubly exponential large coefficients. For example, consider
$x^{(n+1)n}\bmod x^{n+1}-x-1=\sum_{0\leq i\leq n}{\binom{n}{i}x^{i}}.$
Here the input size is $O(\log n)$, however, the coefficient of $x^{n/2}$ of the remainder is of size $2^{n-O(\log n)}$, which is doubly exponential in the input size. This rules out any approach that computes the remainder explicitly. Nevertheless, a solution to the divisibility testing problem has to decide whether the remainder is zero or not. Another challenge arises from the inefficiency of evaluating polynomials: the evaluation of a sparse polynomial $f\in\mathbb{Z}[x]$ on an algebraic number $a$ may have an exponential bit complexity, unless $\left|a\right|=1$ or $a=0$. Thus, natural approaches to the divisibility problem are not effective when considering sparse polynomials of high degree. Prior to our work, the best upper bounds on the size of the coefficients of the quotient polynomial, when $g$ exactly divides $f$, were exponentially large in the degree of $f$, see [Mah62], [BG06, Chapter 1.6] and [Mig12, Proposition 4.9]. Thus, it was previously unknown whether coefficients of the quotient polynomial have “reasonable” representation size, even when the divisor polynomial has a bounded number of terms and the remainder is zero.
Some of the obstacles described above disappear when working over finite fields, such as having too large coefficients, or not being able to evaluate efficiently. However, there are reasons to believe that divisibility testing is a hard problem over finite fields as well. Bi, Cheng, and Rojas proved that with respect to BPP-reductions, deciding whether two sparse polynomials have non-trivial common divisor or whether a sparse polynomial is square-free is NP-Hard. [BCR16]. Other related problems such as computing the number of common roots of $n$ sparse polynomials $f_{1},\ldots,f_{n}\in\mathbb{F}_{p}[x]$, of degrees $\exp(n)$, were shown to be #$\#P$ complete [Qui86, vzGKS93].
This work:
We make significant progress toward solving 2 in the exact case and improve the bound given in Equation (1.1) on the running time of the exact division algorithm.
Specifically, we prove that, when $g$ exactly divides $f$ it holds that
$\log\|f/g\|_{\infty}\leq\log\|f\|_{\infty}+\tilde{O}\left(\|g\|_{0}\right)\cdot\log\deg f.$
For example, our result shows that when $g$ has a bounded number of terms, the coefficients of the quotients are at most polynomial in $\deg f$, thus yielding an exponential improvement over classical results [Mah62], [BG06, Chapter 1.6] and [Mig74].
An implication of this result is that, in the exact case, the running time of the long division algorithm is
$\tilde{O}(\|f/g\|_{0}\cdot\|g\|_{0}\cdot(\log\|f\|_{\infty}+\log\|g\|_{\infty}+\|g\|_{0}\cdot\log\deg f)),$ (1.2)
where earlier results could only prove an upper bound of the form $\Omega(\|f/g\|_{0}\cdot\|g\|_{0}\cdot\deg f)$ (see Equation (1.1)) [DC09].
@@
We next discuss more related work and then present our results.
1.3 Related Work
1.3.1 Bounds on the Coefficients of the Quotient Polynomial
Let $f,g,h\in\mathbb{F}[x]$ such that $f=gh$. As mentioned in [GGPdC21, Lemma 2.12], a simple induction on the long division algorithm yields the following bound on the coefficients of $h$:
This bounds the representation size of the coefficients of $h$ by a linear function of $\left\|{h}\right\|_{0}$. However, when the quotient $h$ is not sparse, this bound becomes exponential in the degree of $h$, which can be exponential in the representation size of $f$ and $g$.
The best bound for this problem is due to Mignotte [Mig74], which is notable for being used in the polynomial factorization algorithm of Lenstra, Lenstra, and Lovász [LLL82].
Theorem 1.4 (Mignotte’s Bound [Mig74, Theorem 2]).
Let $f,g,h\in\mathbb{Q}[x]$ such that $f=gh$. Then,
$\left\|{h}\right\|_{1}\leq 2^{\deg{f}}\frac{\left\|{f}\right\|_{2}}{\left\|{g}\right\|_{1}}.$
Since in our case $\left\|{g}\right\|_{1}$ is relatively small, it is also worth noting another version of this bound which states that
$\left\|{h}\right\|_{1}\leq 2^{\deg{h}}\left\|{f}\right\|_{2}$ (1.5)
for any factor $h$ of $f$. Discussions about this classical inequality can be found in [Mig88, Section 2] and [Mig12, Chapter 4]. These bounds are also proven to be somewhat tight; for any $d\in\mathbb{N}$ there exists $h\in\mathbb{Z}[x]$ of degree $d$ and $f\in\mathbb{Z}[x]$ such that
Another classical bound is known as Gel’fond’s Lemma (see [Mah62] and [BG06, Lemma 1.6.11]).
Theorem 1.6 ( Gel’fond’s Lemma).
Let $f,g,h\in\mathbb{Q}[x]$ such that $f=gh$. Then,
However, none of the above bounds is useful for divisibility testing when the degrees of $f,g$ and $h$ are exponentially large.
1.3.2 Exact Division of Sparse Polynomials
Monagan and Pearce [MP11, MP07] gave an algorithm for computing the quotient and remainder that performs $O(\|f/g\|_{0}\cdot\|g\|_{0})$ many ring operations. This algorithm ignores the bit complexity of the representation as, by the examples given above, the coefficients of the remainder polynomial can be doubly exponentially large in the input size. Their algorithm runs well when the result has small representation size.
In [GGPdC21, GGPdCR22] Giorgi, Grenet, Perret du Cray and Roche studied the exact divisibility problem over both finite fields and integers. They proposed randomized algorithms that run in quasi-linear time in the input size and representation size of the quotient polynomial. However, prior to our work, over the integers (or large finite fields), the best bounds on the coefficient size of the quotient were exponential in the input size. Hence, the only provable bound on the running time, in terms of the input and output size, was known to be quadratic. Moreover, prior to our work no deterministic quasi-linear time algorithm in the representation size of the quotient polynomial was known.
[…]
1.3.3 Testing Divisibility of Sparse Polynomials
[…]
A notable related result is due to Filaseta and Schinzel [FS04], who gave a polynomial time algorithm for deciding whether a sparse polynomial over the integers, with a bounded number of non-zero terms (i.e. $O(1)$ many terms), has a cyclotomic factor. In the general case, when the sparsity is not bounded, this problem was shown to be NP-complete [CTV10].
In [FGS08] Filaseta, Granville, and Schinzel showed that if $f$ and $g$ have bounded sparsities, bounded heights, and at least one of them is cyclotomic-free, then testing whether $g$ divides $f$ can be solved in linear time by computing their greatest common divisor.
A recent work by Giorgi, Grenet, and Perret du Cray [GGPdC21] shows that there exists a polynomial time algorithm for divisibility testing of sparse polynomials, in several special cases. Roughly, they proved that when there is a certain gap between the highest and second highest exponents in the divisor polynomial $g$ then the sparsity of the quotient polynomial can be bounded (which enables them to check divisibility using long division). Another special case that they handle is when $g=g_{0}+x^{k}g_{1}$ where $k=\Omega(\deg{f}-\deg{g})$ and $\deg{g_{0}}$ (or $\deg{g_{1}}$) is poly-logarithmic in the input size. A special case of their result applies to testing divisibility by trinomials; specifically, it implies that there exists a polynomial time algorithm that decides whether $g\mid f$ when $g$ is a trinomial and $\deg{f}=O(\deg{g})$. We obtain such a result for all cyclotomic-free trinomials $g$ and sparse polynomials $f$, regardless of their degrees.
1.4 Our Results
Our first main result gives a new bound on the $\ell_{2}$-norm of a quotient of two sparse polynomials over $\mathbb{C}$, thus improving the classical bounds of Mignotte [Mig74] and Gel’fond [Mah62] on the height of the quotient polynomial when $g$ is sparse.
Theorem 1.8 (Restatement of Theorem 3.10).
Let $f,g,h\in\mathbb{C}[x]$ such that $f=gh$ and $g$ is monic. Then,
$\left\|{h}\right\|_{2}\leq\left\|{f}\right\|_{1}\tilde{O}\left(\left\|{g}\right\|_{0}^{3}\deg^{2}{f}\right)^{\left\|{g}\right\|_{0}-1}.$
As opposed to the bounds given in Equations (1.3), (1.5) and (1.7), our bound on the coefficients of $h$ is not exponential in the degree of $g$, but rather in its sparsity, which for sparse polynomials may be exponentially smaller. Even more surprisingly, the bound in (1.9) does not depend on the height of $g$, in spite of the “feedback process” in the long division algorithm. We also get the following bound as an immediate corollary of Theorem 1.8.
$\log_{2}{\left\|{h}\right\|_{\infty}}\leq\log_{2}{\left\|{f}\right\|_{\infty}}+\tilde{O}(\|g\|_{0})\cdot\log_{2}{\deg{f}}.$ (1.9)
Thus, when $\|g\|_{0}$ is small this gives an exponential improvement over the bounds of Mignotte and Gel’fond (Theorems 1.41.6). In addition, Theorem 1.8 gives an exponential separation between exact and non-exact division since $\log_{2}{\left\|{h}\right\|_{\infty}}$ can be as large as $\Omega(\deg{f}\cdot\log_{2}{\left\|{g}\right\|_{\infty}})$ when allowing non-zero remainder.
As a corollary of Theorem 1.8 we obtain the following guarantee on the performance of the long division algorithm in the exact case.
Theorem 1.10.
When the division is exact, the time complexity of the long division algorithm for polynomials $f,g\in\mathbb{Z}[x]$ is
$\tilde{O}\left(\left\|{f/g}\right\|_{0}\cdot\left\|{g}\right\|_{0}\cdot\left(\log{\left\|{g}\right\|_{\infty}}+\log{\left\|{f}\right\|_{\infty}}+\left\|{g}\right\|_{0}\log{\deg{f}}\right)\right).$
[…]
@@ 1.4 Our Results
Combining the bound in Theorem 1.8 with Theorem 1.3 of [GGPdCR22] we obtain the following corollary.
@@
$\tilde{O}\left((\left\|{f}\right\|_{0}+\left\|{g}\right\|_{0}+\left\|{f/g}\right\|_{0})\cdot\left(\log{\left\|{g}\right\|_{\infty}}+\log{\left\|{f}\right\|_{\infty}}+\left\|{g}\right\|_{0}\log{\deg{f}}\right)\right).$
Next, we prove new bounds on the sparsity of the quotient of two polynomials over $\mathbb{Z}$ when $g$ is a cyclotomic-free binomial or a cyclotomic-free trinomial.
Theorem 1.12 (Restatement of Theorem 4.6).
Let $f,g,h\in\mathbb{Z}[x]$ such that $f=gh$. If $g$ is cyclotomic-free, and contains two terms then the sparsity of $h$ is at most
$O\left(\left\|{f}\right\|_{0}(\log{\left\|{f}\right\|_{0}}+\log{\left\|{f}\right\|_{\infty}})\right).$
Theorem 1.13 (Restatement of Theorem 4.17).
Let $f,g,h\in\mathbb{Z}[x]$ such that $f=gh$. If $g$ is cyclotomic-free and contains three terms, then the sparsity of $h$ is at most
$O\left(\left\|{f}\right\|_{0}\cdot\text{size}\left(f\right)^{2}\cdot\log^{6}{\deg{g}}\right).$
The above bounds give rise to explicit algorithms for testing the divisibility of polynomials over the integers by a binomial and by a cyclotomic-free trinomial. In what follows, $n$ will denote the size of the input, i.e., $\text{size}\left(f\right)+\text{size}\left(g\right)$.
Theorem 1.14.
Let $f,g\in\mathbb{Z}[x]$ be two polynomials given in the sparse representation, such that $g$ contains two terms. Then, there exists an algorithm that decides whether $g\mid f$ (and if so, computes the quotient $f/g$) in $\tilde{O}(n^{2})$ time.
Theorem 1.15.
Let $f,g\in\mathbb{Z}[x]$ be two polynomials given in the sparse representation, such that $g$ is cyclotomic-free and contains three terms. Then, there exists an algorithm that decides whether $g\mid f$ (and if so, computes the quotient $f/g$) in $\tilde{O}(n^{10})$ time.
Theorem 1.15 gives a positive answer to 1 for the case of trinomials over the integers.
Finally, we address the problem of polynomial divisibility over finite fields and show that testing divisibility by a binomial can be solved efficiently.
Theorem 1.16.
Let $f,g\in\mathbb{F}_{p}[x]$ be two polynomials given in the sparse representation, such that $g$ contains two terms. Then, there exists an algorithm that decides whether $g\mid f$ in $\tilde{O}(n^{2})$ time.
Binomials are a special case of a large class of polynomials, for which our algorithm also applies.
Remark 1.17.
[…]
As our last result, we present a new algorithm for divisibility testing by a pentanomial when $\deg{f}=\tilde{O}(\deg{g})$.
Theorem 1.18.
[…]

### v4

# A New Bound on Cofactors of Sparse Polynomials

@@ Abstract
We prove that for polynomials $f,g,h\in\mathbb{Z}[x]$ satisfying $f=gh$ and $f(0)\neq 0$, the $\ell_{2}$-norm of the cofactor $h$ is bounded by
$\|h\|_{2}\leq\|f\|_{1}\cdot\left(\widetilde{O}\left(\|g\|_{0}^{3}\frac{\deg(f)^{2}}{\sqrt{\deg(g)}}\right)\right)^{\|g\|_{0}-1},$
where $\|g\|_{0}$ is the number of nonzero coefficients of $g$ (its sparsity). We also obtain similar results for polynomials over $\mathbb{C}$ [^1] These bounds improve upon the conference version of this paper [NS24].
This result significantly improves upon previously known exponential bounds (in $\deg(f)$) for general polynomials. It further implies that, under exact division, the polynomial division algorithm runs in quasi-linear time with respect to the input size and the number of terms in the quotient $h$. This resolves a long-standing open problem concerning the exact divisibility of sparse polynomials.
In particular, our result demonstrates a quadratic separation between the runtime (and representation size) of exact and non-exact divisibility by sparse polynomials. Notably, prior to our work, it was not even known whether the representation size of the quotient polynomial could be bounded by a sub-quadratic function of its number of terms, specifically of $\deg(f)$.
@@ 1 Introduction
Let $f,g,h\in\mathbb{C}[x]$ be polynomials such that $f=gh$. Classical results of Gel’fond, Mahler and Mignotte provide upper bounds on the height of $h$ ($\ell_{\infty}$ norm of its vector of coefficients) given information on the coefficients of $g$ and $f$ [Gel60, Mah62, Mig74, Mig88]. These results were used by Gel’fond in the study of transcendence and are widely used in computer algebra to prove upper bounds on the running time of algorithms, such as in the polynomial factorization algorithm of Lenstra, Lenstra, and Lovász [LLL82].
The first of these bounds was proved by Gelf’ond [Gel60] (see also [Mah62] and [BG06, Lemma 1.6.11]).
Theorem 1.1 ( Gel’fond’s Lemma).
Let $f,g,h\in\mathbb{C}[x]$ such that $f=gh$. Then,
Improved bounds were provided by Mahler [Mah62] and consequently by Mignotte [Mig74].
Theorem 1.3 (Mignotte’s Bound [Mig74, Theorem 2]).
Let $f,g,h\in\mathbb{Z}[x]$ such that $f=gh$. Then,
$\left\|{h}\right\|_{1}\leq 2^{\deg{h}}\left\|{f}\right\|_{2}$ (1.4)
Discussions about this classical inequality can be found in [Mig88, Section 2] and [Mig12, Chapter 4]. This bound is also known to be quite tight; for any $d\in\mathbb{N}$ there exists $h\in\mathbb{Z}[x]$ of degree $d$ and $f\in\mathbb{Z}[x]$ such that
When $\|h\|_{0}$ is guaranteed to be small relative to its degree, tighter bounds can be obtained. For instance, Giorgi, Grenet and Perret du Cray established the following bound by a straightforward induction on the polynomial division algorithm [GGPdC21].
Lemma 1.5 (Lemma 2.12 in [GGPdC21]).
Let $f,g,h\in\mathbb{Z}[x$] satisfy $f=gh$. Then
While the bound given in (1.4) is close to being tight, in this paper we ask whether one can get an improvement if the factor $g$ is sparse. For example, Khovanskiĭ’s theory asserts that when studying real roots of polynomial equations, results analogous to Bézout’s theorem can be obtained, where the number of monomials in the equations takes the place of the degree in Bézout’s theorem [Kho91].
Besides being a self-motivated question, bounding the norm of cofactors of sparse polynomials have important applications in computer algebra which we discuss next.
1.1 The Sparse Representation
@@
where $\left\|{f}\right\|_{0}$ and $\left\|{f}\right\|_{\infty}$ are the number of non-zero coefficients in $f$, and the maximal coefficient of $f$ (in absolute value), respectively.222 As we think of $\left\|{f}\right\|_{0}$ as being considerably smaller than $\log_{2}{\left\|{f}\right\|_{\infty}}$ and $\log_{2}{\deg{f}}$, we do not pay attention to the case where bit lengths of coefficients vary. This makes the sparse representation especially useful for applications involving polynomials with high degrees where most of the coefficients are zero. In particular, it is used by computer algebra systems and libraries such as Maple [MP13], Mathematica [Wol99], Sage [S+08], and Singular [Sch03].
@@
One of the long-standing open problems concerns exact divisibility of sparse polynomials. There are two versions to this problem. The decision problem asks, given two sparse polynomials $f$ and $g$, to decide whether $g$ divides $f$. The search problem, or the exact division problem, asks to compute the quotient polynomial $f/g$ when it is known that $g$ divides $f$.
The two problems are closely related but also differ in what we can hope to achieve. In the decision problem, the output is a single bit, allowing us to analyze the complexity in terms of the input size. Conversely, in the search (exact division) problem, the output size depends on the number of monomials in $f/g$, which may require an exponential number of bits (compared to the size of the input) to represent.
There are two quantities that affect the complexity of the exact division problem. The first is the number of terms of the quotient $f/g$. It is easy to find examples in which $f/g$ has exponentially many monomials. The second quantity is the height. Prior to this work the only bounds on the height were exponential in $\deg{f}$, as in Equations (1.2),(1.4) and(1.6) (see also [Mig12]). In particular, Davenport and Carette note that even in the case where $g$ exactly divides $f$, we need to perform $\Omega(\|f/g\|_{0}\cdot\|g\|_{0})$ operations on coefficients, and the number of bits required to represent a coefficient may be as large as $\deg f$ [DC09, Section III.B]. Therefore, in term of the input size and $\|f/g\|_{0}$ the only provable bounds on the running time of the exact division algorithm were of the form
@@
Following [DC09], Roche posed the next problem [Roc18]. [^3] We use the convention, $\widetilde{O}(t)=O\left(t\cdot\log^{O(1)}t\right)$ , and when no basis is mentioned then the log is taken in the natural base $e$.
Open Problem 1 ([Roc18, Open Problem 3]).
Given two sparse polynomials $f,g\in R[x]$, develop an algorithm which determines whether $g$ divides $f$ exactly, using $\widetilde{O}(T\cdot\log{\deg{f}})$ ring and bit operations, [^4] In the study of algorithms over rings, we distinguish between operations on ring elements, which may depend on the ring’s representation, and operations on bit strings. For instance, ring operations include addition, multiplication, and division of coefficients, while bit operations correspond to addition and subtraction of exponent vectors, typically represented as bit strings. In practical scenarios involving domains like $\mathbb{Z}$ or $\mathbb{C}$, the computational cost of each ring operation on a computer is quasi-linear in the number of bits used to represent the relevant elements. where $T$ is the total number of terms of $f$, $g$, $(f\text{ quo }g)$ and $(f\text{ rem }g)$ (the quotient and remainder).
@@
Note that 1 asks for an algorithm for the decision problem, whose running time may be exponential in the input size (as in the search problem).
1.2 Our Results
Our main result gives a new bound on the $\ell_{2}$-norm of $f/g$, for $f,g\in\mathbb{C}[x]$ such that $g$ divides $f$. This result improves the classical bounds of Gel’fond [Mah62] and Mignotte [Mig74] when the factor $g$ is sparse. In what follows we assume without loss of generality that $f(0)\neq 0$ as we can always remove powers of $x$ that divide the polynomials without affecting their coefficients. To slightly simplify the expression, we assume that $f(0)\neq 0$. If this is not the case, then $\deg{g}$ should be replaced by $\deg{g}-\text{ord}_{0}{g}$ throughout.
Theorem 1.8.
Let $f,g,h\in\mathbb{C}[x]$ such that $f=gh$ and $f(0)\neq 0$. Denote $L(n)=n\log{n}$. Let
$\text{M}=\max\left\{\begin{array}[]{c}\min\{\left|\text{LC}(g)\right|,\left|\text{TC}(g)\right|\}\cdot\left(\frac{\|g\|_{0}}{2e}\right)^{\frac{\|g\|_{0}-1}{2}}\cdot\left(\frac{\deg{g}}{e}\right)^{\frac{\left\|{g}\right\|_{0}}{2}}\\ \max\{\left|\text{LC}({g})\right|,\left|\text{TC}({g})\right|\}\cdot\deg{g}\end{array}\right\}.$
Then,
$\left\|{h}\right\|_{2}\leq\frac{\sqrt{2}\left\|{f}\right\|_{1}}{\text{M}{}}\left({2\left\|{g}\right\|_{0}\cdot L\left(12\left\|{g}\right\|_{0}\deg{g}+2\deg{f}\right)^{2}}\right)^{\left\|{g}\right\|_{0}-1}.$
Remark 1.9.
We note that we must have a term that inversely depend on $|\text{LC}(g)|$ or $|\text{TC}(g)|$ since if $hg=f$ then for every constant $c$ we also have $(c\cdot h)\cdot(g/c)=f$. This scales $\|h\|_{2}$ by a factor of $c$.
The following is an immediate corollary.
Corollary 1.10.
Let
  • $f,g,h\in\mathbb{Z}[x]$ such that $f=gh$ and $f(0)\neq 0$, or
  • $f,g,h\in\mathbb{C}[x]$ such that $f=gh$, $f(0)\neq 0$, and $|\text{LC}(g)|,|\text{TC}(g)|\geq 1$.
Then, it holds that
$\left\|{h}\right\|_{2}\leq\sqrt{\frac{\left\|{g}\right\|_{0}}{\deg g}}\cdot\|f\|_{1}\cdot\left(\widetilde{O}\left(\frac{\left\|{g}\right\|_{0}^{2.5}\cdot\mathrm{deg}^{2}{g}+\sqrt{\|g\|_{0}}\cdot\mathrm{deg}^{2}{f}}{\sqrt{\deg{g}}}\right)^{2}\right)^{\left\|{g}\right\|_{0}-1}.$
In terms of bit representation size we have
$\log_{2}{\left\|{h}\right\|_{\infty}}\leq\log_{2}{\left\|{h}\right\|_{2}}\leq\log_{2}{\left\|{f}\right\|_{\infty}}+O\left(\|g\|_{0}\cdot\log_{2}{\deg{f}}\right).$ (1.11)
Corollary 1.12.
Let $f,g\in\mathbb{C}[x]$ such that $f=gh$ and $g$ is monic. Then it holds that
$\left\|{h}\right\|_{2}\leq\frac{\|f\|_{1}}{\deg{g}}\cdot\widetilde{O}\left(\left\|{g}\right\|_{0}^{3}\mathrm{deg}^{2}{f}\right)^{\left\|{g}\right\|_{0}-1}.$
Unlike the bounds in Equations (1.2), (1.4) and (1.6), our bound on the height of $h$ is not exponential in the degree or number of coefficients of $h$, but rather in the sparsity of the factor $g$, which for sparse polynomials may be exponentially smaller. Thus, when $\|g\|_{0}$ is small, this gives an exponential improvement over the bounds of Gel’fond and Mignotte (Theorems 1.11.3). In addition, Theorem 1.8 gives an exponential separation between exact and non-exact division since $\log_{2}{\left\|{h}\right\|_{\infty}}$ can be as large as $\Omega(\deg{f}\cdot\log_{2}{\left\|{g}\right\|_{\infty}})$ when allowing non-zero remainder.
Theorem 1.8 also provides the following guarantee regarding the performance of the polynomial division algorithm in the exact case.
Theorem 1.13.
When the division is exact, the time complexity of the polynomial division algorithm for polynomials $f,g\in\mathbb{Z}[x]$ is
$\widetilde{O}\left(\left\|{f/g}\right\|_{0}\cdot\left\|{g}\right\|_{0}\cdot\left(\log{\left\|{g}\right\|_{\infty}}+\log{\left\|{f}\right\|_{\infty}}+\left\|{g}\right\|_{0}\log{\deg{f}}\right)\right).$
@@ 1.2 Our Results
Finally, combining the bound in Theorem 1.8 with Theorem 1.3 of [GGPdCR22] we obtain the following result.
@@
$\widetilde{O}\left((\left\|{f}\right\|_{0}+\left\|{g}\right\|_{0}+\left\|{f/g}\right\|_{0})\cdot\left(\log{\left\|{g}\right\|_{\infty}}+\log{\left\|{f}\right\|_{\infty}}+\left\|{g}\right\|_{0}\log{\deg{f}}\right)\right).$